# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v2

@@ Abstract
We introduce an algorithm for the solution of a large system of radial Schrödinger equations for scattering states. The system of differential equations is approximated as an ordinary linear nonhomogeneous system using the finite difference method. Dirichlet boundary conditions are imposed at the origin and at an arbitrary large radius. The $K$-matrix for physical energies is calculated from the numerical solutions of the system by comparison to the analytical real solutions at large distances. Scattering poles and amplitudes are calculated from the physical $K$-matrix.
@@ 1 Introduction
The Schrödinger equation applies to scattering states just as well as to bound states. In fact, the Schrödinger equation provides the foundation of many approximate analytical methods to calculate scattering poles and amplitudes using nonrelativistic quantum mechanics. But the advantages of calculating scattering poles and amplitudes directly from the Schrödinger equation are often overlooked.
@@
The most obvious benefit of calculating scattering states directly from the Schrödinger equation is that it does not require any secondary approximation that may deteriorate the accuracy of the results. The scattering wavefunctions are completely determined by the Schrödinger equation, which in turn is completely determined by the potential and the masses of the degrees of freedom. Therefore, the scattering poles and amplitudes obtained in this way are just as good as the model or framework in which the Schrödinger equation is defined. That is most advantageous for those cases where resonant line shapes overlap with one another or with the energy thresholds where scattering channels open up. There is no need for a special treatment of esoteric structures like dips and cusps in the scattering amplitudes, since they are automatically included in the Schrödinger equation.
@@
A system of Schrödinger equations can be used to describe the quantum mechanical scattering of particles in systems where different final states are possible after a collision at nonrelativistic center-of-mass energies. One notable example is nucleon-nucleon scattering, where inelastic effects start to play a role at relatively low scattering momenta ($NN\to NN,N\Delta,NN^{\prime},\Delta\Delta,\dots$).
@@
In this paper, we introduce a simple yet effective algorithm for calculating scattering poles and amplitudes directly from a system of radial Schrödinger equations. Our method is to solve the system using a simple finite-difference method (see Chapter 3 of Reference [1]). Other common methods include propagation algorithms [2, 3] or basis expansions [4, 5] (see Reference [1] for a review). The crudeness of the finite-difference method is offset by its efficiency, which makes it the gold standard for numerically solving a differential equation with minimal effort.
@@
In phenomenological applications, there are many situations in which one has to solve a system of Schrödinger equations several times to obtain a definite prediction. For instance, one may need to fit the parameters of some potential model, or one may want to estimate the error on the calculated spectrum from the uncertainty of the potential parameters using a bootstrap method. In those cases, it is desirable to have an algorithm that solves the system of Scrhödinger equations as efficiently as possible. The extreme efficiency of the finite-difference method allows to solve systems with tens of coupled Schrödinger equations [^2] Such as the equations governing certain strongly interacting systems in the Born-Oppenheimer approximation for quantum chromodynamics; see, for instance, Ref. [6]. using only modest computing power and time. That is the main purpose and novelty of the SPARSE algorithm.
@@
The rest of the paper is structured as follows. In Section 2, we derive the system of radial Schrödinger equations describing an inelastic two-body scattering process in a partial wave with definite total-angular-momentum quantum numbers. In Section 3, we reduce the system of radial Schrödinger equations down to a nonhomogeneous linear system. In Section 4, we calculate the reactance matrix $K$ for the scattering process by comparing the numerical wavefunctions to the analytic ones for scattering states with real boundary conditions. In Section 5, we review the calculation of scattering poles and amplitudes from the $K$-matrix. In Section 6, we showcase the SPARSE algorithm using a simple example that serves also as a tutorial. Finally, we summarize these results in Section 7.
2 Radial Schrödinger equation
@@
A two-body elastic or inelastic scattering process at nonrelativistic energies can be described by a system of coupled Schrödinger equations. Using matrix notation, the system can be written concisely as
@@
where $\bm{r}_{1}$, $\bm{r}_{2}$ are the particles’ coordinates , $\bm{p}_{1}=-i\bm{\nabla}_{\bm{r}_{1}}$, $\bm{p}_{2}=-i\bm{\nabla}_{\bm{r}_{2}}$ are their conjugate momenta, $m_{1}$ and $m_{2}$ are diagonal mass matrices, $V(\bm{r}_{1}-\bm{r}_{2})$ is a spin-dependent potential matrix, $E$ is the total energy, and $\psi(\bm{r}_{1},\bm{r}_{2})$ is a vector wavefunction. Note that we use natural units, $\hbar=c=1$. Each Schrödinger equation in the system governs an elastic scattering process in a given channel. Each diagonal element of the mass matrices $m_{1}$ and $m_{2}$ contains the mass of the particles 1 and 2 in the corresponding channel. The couplings between the different Schrödinger equations introduce inelastic effects. In general, Equation (1) describes interacting particles with spin. Each component of the vector wavefunction $\psi(\bm{r}_{1},\bm{r}_{2})$ contains the spin states of the scattering particles in the corresponding channel. Each element of the potential matrix $V(\bm{r}_{1}-\bm{r}_{2})$ is an operator acting on the spin states.
@@
where $\overline{m}_{1}$ and $\overline{m}_{2}$ are the average of the masses in the diagonal matrices $m_{1}$ and $m_{2}$. We also introduce the relative momentum
@@
where $E$ now represents the energy in the center-of-momentum frame where $\bm{P}=0$. Note that the separability condition in Equation (8) is satisfied exactly if and only if the two mass matrices are linearly dependent, $m_{1}\propto m_{2}$. If $m_{1}$ and $m_{2}$ are linearly independent, Equation (8) cannot be satisfied exactly but Equation (6) may be approximately separable if
@@
where the double vertical lines denote the matrix norm.
We now wish to further reduce Equation (10) to a system of differential equations in the distance $r=\lvert\bm{r}\rvert$. Let us label the different Schrödinger equations in the system by an integer $a$. The Schrödinger equation in channel $a$ reads
@@
where $\psi_{a}(\bm{r})$ is the wavefunction in channel $a$ and $V_{ab}(\bm{r})$ are potential matrix elements. Recall that $\psi_{a}(\bm{r})$ contains the spin states of the scattering particles and $V_{ab}(\bm{r})$ is an operator acting on the spin states. Let $J_{1,a}$ and $J_{2,a}$ be the spins of the particles 1 and 2 in the channel $a$. We can expand the wavefunction $\psi_{a}(\bm{r})$ as
@@
where $\psi_{a,s,l}^{\sigma,m}(r)$ are radial wavefunctions, $Y_{l}^{m}(\theta,\phi)$ are spherical harmonics, and $\xi_{s}^{\sigma}$ are irreducible representations from the decomposition of the direct product of the spin states $\chi_{J_{1,a}}^{M_{1,a}}$ and $\chi_{J_{2,a}}^{M_{2,a}}$ of the particles 1 and 2,
@@
where the term with angle brackets is a Clebsch-Gordan coefficient. The sum over the total spin quantum number $s$ ranges from $\lvert J_{1,a}-J_{2,a}\rvert$ to $J_{1,a}+J_{2,a}$, and for each $s$ the sum over the third component $\sigma$ ranges from s-s to +s+s. The sum over the orbital-angular-momentum quantum number $l$ ranges from 0 to infinity, and for each $l$ the sum over the third component $m$ ranges from l-l to +l+l. The radial wavefunctions in Equation (13) are labeled by quantum numbers $\sigma$ and $m$ for the third components of the spin and orbital angular momentum, respectively. It is more convenient to rewrite Equation (13) as an expansion in radial wavefunctions labeled by quantum numbers $J$ and $M$ for the total angular momentum,
@@
Once the above procedure has been applied to each channel $a$, one obtains infinitely many systems of radial Schrödinger equations labeled by conserved quantum numbers $J$ and $M$ for the total angular momentum and its third component. Using matrix notation, the system for given quantum numbers $J$ and $M$ can be written concisely as
$\biggl[\frac{1}{2\mu}\biggl(-\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}-\frac{2}{r}\frac{\mathrm{d}\hphantom{r}}{\mathrm{d}r}+\frac{L(L+1)}{r^{2}}\biggr)+V^{J}(r)-E^{J}\biggr]\psi^{JM}(r)=0,$ (16)
@@
where $L$ is a diagonal orbital-angular-momentum matrix, $V^{J}(r)$ is a radial potential matrix, and $\psi^{JM}(r)$ is a vector wavefunction. Note that the wavefunctions depend on both $J$ and $M$, but the potential and the energy depend only on $J$. The radial Schrödinger equations in this system are labeled by triplets $(a,l,s)$ that identify spin-orbital channels of the scattering process. Since $s$ and $l$ are constrained by the triangle relations $\lvert J_{1,a}-J_{2,a}\rvert\leq s\leq J_{1,a}+J_{2,a}$ and $\lvert J-s\rvert\leq l\leq J+s$, the system in Equation (16) contains only a finite number of radial Schrödinger equations for any value of $J$. To simplify the notation, we replace the labels $(a,l,s)$ for the spin-orbital channels by a single integer $i$. Then the radial potential matrix elements can be expressed as
@@
$V_{ij}^{J}(r)=\sum_{\sigma_{i},m_{i}}\sum_{\sigma_{j},m_{j}}\langle s_{i}l_{i}\sigma_{i}m_{i}|s_{i}l_{i}JM\rangle\langle s_{j}l_{j}\sigma_{j}m_{j}|s_{j}l_{j}JM\rangle\\ \times\int\mathrm{d}\Omega\,Y_{l_{i}}^{m_{i}}(\theta,\phi)^{\ast}Y_{l_{j}}^{m_{j}}(\theta,\phi)\xi_{s_{i}}^{\sigma_{i}{\dagger}}V_{ab}(\bm{r})\xi_{s_{j}}^{\sigma_{j}},$ (17)
@@
where $(a,s_{i},l_{i})\to i$, $(b,s_{j},l_{j})\to j$, and the total spin states $\xi_{s_{i}}^{\sigma_{i}}$ and $\xi_{s_{j}}^{\sigma_{j}}$ can be expanded in terms of the spin states $\chi_{J_{1,i}}^{M_{1,i}},\chi_{J_{2,i}}^{M_{2,i}}$ and $\chi_{J_{1,j}}^{M_{1,j}},\chi_{J_{2,j}}^{M_{2,j}}$ of particles 1, 2 in the channel $i$ and $j$ using Equation (14). Note that the right side of Equation (17) does not depend on $M$ by virtue of the Wigner-Eckart theorem.
@@
There is a somewhat common misunderstanding that the reduction of a three-dimensional Schrödinger equation to a radial form can be performed only for a spherically-symmetric potential that does not depend on the angles $\theta$ and $\phi$. However, in our derivation of Equations (16) and (17) we have made no assumption on the potential $V(\bm{r})$ other than it conserves angular momentum. In the trivial case where the potential is independent of spin, then conservation of angular momentum does indeed require that it is also spherically symmetric. If the potential does depend on spin, then it can also depend on the angles $\theta$ and $\phi$ but Equations (16) and (17) would still be valid. A brute-force method to determine if any given potential $V(\bm{r})$ conserves angular momentum is to evaluate the right side of Equation (17) with $J$ and $M$ in one of the two Clebsch-Gordan coefficients replaced by $J^{\prime}$ and $M^{\prime}$. The potential $V(\bm{r})$ conserves angular momentum if and only if the result is zero for $J^{\prime}\neq J$ or $M^{\prime}\neq M$ and it is independent of $M$ when $J^{\prime}=J$ and $M^{\prime}=M$.
@@
Equation (16) describes a scattering process taking place in a partial wave labeled by quantum numbers $J$ and $M$. Because of conservation of angular momentum, each partial wave is decoupled from the others. SPARSE solves the scattering problem only for one individual partial wave, leaving the task of summing partial waves to the user. Hence, from now on we will drop all superscripts for the conserved quantum numbers $J$ and $M$.
@@
Finally, we can simplify the derivatives inside Equation (16) as follows. Let us observe that

### v3

@@ Abstract
We introduce an algorithm for the solution of a system of radial Schrödinger equations describing the inelastic scattering of particles with spin in a partial wave with definite total angular momentum. The system of differential equations is approximated as an ordinary linear nonhomogeneous system using the finite difference method. Dirichlet boundary conditions are imposed at the origin and at an arbitrary large radius. The $K$-matrix for physical energies is calculated from the numerical solutions of the system by comparison to the analytical real solutions at large distances. Scattering poles and amplitudes are calculated from the physical $K$-matrix.
@@ 1 Introduction
The Schrödinger equation applies to scattering states as well as to bound states. In fact, the Schrödinger equation provides the foundation of many approximate analytical methods to calculate scattering poles and amplitudes using nonrelativistic quantum mechanics. But the advantages of calculating scattering poles and amplitudes directly from the Schrödinger equation are often overlooked.
@@
The most obvious benefit of calculating scattering states directly from the Schrödinger equation is that it does not require any secondary approximation that may limit the validity of the results. The scattering wavefunctions are completely determined by the Schrödinger equation, which in turn is completely determined by the potential and the masses of the degrees of freedom. Therefore, the scattering poles and amplitudes obtained in this way are as good as the model or framework in which the Schrödinger equation is defined. That is most advantageous for those cases where resonant line shapes overlap with one another or with the energy thresholds where scattering channels open up. There is no need for a special treatment of esoteric structures like dips and cusps in the scattering amplitudes, since they are automatically included in the Schrödinger equation.
@@
A system of Schrödinger equations can be used to describe the quantum mechanical scattering of particles with spin in systems where different final states are possible after a collision at nonrelativistic center-of-mass energies. One notable example is nucleon-nucleon scattering, where inelastic effects start to play a role at relatively low scattering momenta ($NN\to NN,N\Delta,NN^{\prime},\Delta\Delta\dots$).
@@
In this paper, we introduce a simple yet effective algorithm for calculating scattering poles and amplitudes directly from a system of radial Schrödinger equations. Our method is to solve the system using a first-order finite-difference method (see Chapter 3 of Reference [undef]). The crudeness of the finite-difference method is offset by its efficiency, which makes it the gold standard for numerically solving a differential equation with minimal computational effort. More sophisticated methods include propagation algorithms [undefa, undefb] and basis expansions [undefc, undefd] (see Reference [undef] for a review).
@@
In phenomenological applications, there are many situations in which one has to solve a system of Schrödinger equations several times to obtain a definite prediction. For instance, one may need to fit the parameters of some potential model, or one may want to estimate the error on the calculated spectrum from the uncertainty of the potential parameters using a bootstrap method. In those cases, it is desirable to have an algorithm that solves the system of Scrhödinger equations as efficiently as possible. That is the main purpose and novelty of the SPARSE algorithm. The extreme efficiency of the finite-difference method allows to solve systems with dozens of coupled Schrödinger equations [^2] Such as the equations governing certain strongly interacting systems in the Born-Oppenheimer approximation for quantum chromodynamics; see, for instance, Ref. [undefe]. using only modest computing power and time.
@@
The rest of the paper is structured as follows. In Section 2, we derive the system of radial Schrödinger equations describing an inelastic two-body scattering process of particles with spin in a partial wave with definite total-angular-momentum quantum numbers. In Section 3, we reduce the system of radial Schrödinger equations to a nonhomogeneous linear system. In Section 4, we calculate the reactance matrix $K$ for the scattering process by comparing the numerical wavefunctions to the analytic ones for scattering states with real boundary conditions. In Section 5, we review the calculation of scattering poles and amplitudes from the $K$-matrix. In Section 6, we showcase the algorithm using a simple example that serves also as a tutorial for SPARSE. Finally, we summarize these results in Section 7.
2 System of radial Schrödinger equations
@@
A two-body elastic or inelastic scattering process at nonrelativistic energies can be described by a system of coupled Schrödinger equations. The system can be written explicitly as
$\sum_{b}\biggl[\delta_{ab}\biggl(\frac{p_{1}^{2}}{2m_{1,a}}+\frac{p_{2}^{2}}{2m_{2,a}}\biggr)+V_{ab}(\bm{r}_{1}-\bm{r}_{2})-\delta_{ab}E\biggr]\psi_{b}(\bm{r}_{1},\bm{r}_{2})=0,$ (1)
where $a$ and $b$ are integers labeling the different Schrödinger equations in the system, also referred to as channels, $\bm{r}_{1}$ and $\bm{r}_{2}$ are the particles’ coordinates, $\bm{p}_{1}=-i\bm{\nabla}_{\bm{r}_{1}}$ and $\bm{p}_{2}=-i\bm{\nabla}_{\bm{r}_{2}}$ are their conjugate momenta, $m_{1,a}$ and $m_{2,a}$ are the masses of the interacting particles in the channel $a$, $V_{ab}(\bm{r}_{1}-\bm{r}_{2})$ is the potential coupling the channel $b$ to the channel $a$, $\psi_{a}(\bm{r}_{1},\bm{r}_{2})$ is the two-body wavefunction in the channel $a$, and $E$ is the total energy. Note that we use natural units, $\hbar=c=1$, and that unitarity requires $V_{ba}(\bm{r}_{1}-\bm{r}_{2})=V_{ab}^{\dagger}(\bm{r}_{1}-\bm{r}_{2})$. In general, Equation (1) describes interacting particles with spin. Each wavefunction $\psi_{a}(\bm{r}_{1},\bm{r}_{2})$ contains the spin states of the interacting particles in the channel $a$ and each potential $V_{ab}(\bm{r}_{1}-\bm{r}_{2})$ is an operator mapping spin vectors in the channel $b$ to spin vectors in the channel $a$. Note also that the system in Equation (1) describes a generally inelastic scattering process where the state of the interacting particles can change during the collision. Therefore, in general one has
$m_{1,a}\neq m_{1,b}\quad\text{and}\quad m_{2,a}\neq m_{2,b}\quad\text{for}\quad a\neq b.$ (2)
Using matrix notation, the system in Equation (1) can be written more concisely as
@@
where $m_{1}$ and $m_{2}$ are diagonal mass matrices, $V(\bm{r}_{1}-\bm{r}_{2})$ is a Hermitian potential matrix, and $\psi(\bm{r}_{1},\bm{r}_{2})$ is a vector wavefunction. The diagonal elements of the mass matrices $m_{1}$ and $m_{2}$ are the masses of the interacting particles in the various Schrödinger equations that make up the system. Each Schrödinger equation, taken in isolation, describes an elastic scattering process. The couplings between the different Schrödinger equations, which are provided by the off-diagonal elements of $V(\bm{r}_{1}-\bm{r}_{2})$, introduce inelastic effects.
@@
where $\overline{m}_{1}$ and $\overline{m}_{2}$ are the average of the masses in the diagonal matrices $m_{1}$ and $m_{2}$, namely,
$\overline{m}_{1}=\frac{\sum_{a}m_{1,a}}{\sum_{a}1}\qquad\text{and}\qquad\overline{m}_{2}=\frac{\sum_{a}m_{2,a}}{\sum_{a}1}.$ (6)
We also introduce the relative momentum
@@
where $E$ now represents the energy in the reference frame where $\bm{P}=0$.
Note that the separability condition in Equation (11) can be satisfied exactly if and only if the two mass matrices are linearly dependent, for it requires $m_{1}\propto m_{2}$. That is always the case for a system of Schrödinger equations describing elastic scattering, where $m_{1}=\overline{m}_{1}I$ and $m_{2}=\overline{m}_{2}I$ with $I$ the identity matrix.
If $m_{1}$ and $m_{2}$ are linearly independent, Equation (11) cannot be satisfied exactly but Equation (9) may be approximately separable if
@@
where the double vertical lines denote the matrix norm. That is typically the case for a system of Schrödinger equations describing inelastic scattering in the nonrelativistic limit. In an inelastic scattering process, the mass difference between the initial and final states gives the amount of kinetic energy released or absorbed during the collision. The nonrelativistic limit corresponds to the assumption that the kinetic energy of each particle is much smaller than its mass. Thus, the mass differences $\lvert m_{1,a}-m_{1,b}\rvert$ and $\lvert m_{2,a}-m_{2,b}\rvert$ with $a\neq b$ can be assumed to be much smaller than the masses $m_{1,a}$ and $m_{2,a}$, respectively. In other words, writing
$\frac{\overline{m}_{1}}{m_{1}}=I+\Delta_{1}\qquad\text{and}\qquad\frac{\overline{m}_{2}}{m_{2}}=I+\Delta_{2},$ (15)
we can assume that the matrices $\Delta_{1}=(\overline{m}_{1}-m_{1})/m_{1}$ and $\Delta_{2}=(\overline{m}_{2}-m_{2})/m_{2}$ are “small”, that is, $\lVert\Delta_{1}\rVert\ll 1$ and $\lVert\Delta_{2}\rVert\ll 1$. Then it follows from Equation (15) and the Cauchy-Schwartz inequality that
$\biggl\lVert\frac{\overline{m}_{1}}{m_{1}}-\frac{\overline{m}_{2}}{m_{2}}\biggr\rVert=\lVert\Delta_{1}-\Delta_{2}\rVert\leq\lVert\Delta_{1}\rVert+\lVert\Delta_{2}\rVert\ll 1.$ (16)
We now wish to further reduce Equation (13) to a system of differential equations in the distance $r=\lvert\bm{r}\rvert$. The Schrödinger equation in the channel $a$ reads
@@
where $\mu_{a}=m_{1,a}m_{2,a}/(m_{1,a}+m_{2,a})$ is the reduced mass in the channel $a$, $V_{ab}(\bm{r})$ is the potential coupling the channel $b$ to the channel $a$, and $\psi_{a}(\bm{r})$ is the wavefunction in the channel $a$. Recall that $\psi_{a}(\bm{r})$ contains the spin states of the interacting particles in the channel $a$, and let $J_{1,a}$ and $J_{2,a}$ be their spins. [^3] Here we refer to the total angular momentum of a particle as its “spin”. This makes it easier to distinguish the total angular momentum of each individual particle from the total angular momentum of the interacting system. We can expand $\psi_{a}(\bm{r})$ as
@@
where $\psi_{a,s,l}^{\sigma,m}(r)$ are radial wavefunctions, $Y_{l}^{m}(\theta,\phi)$ are spherical harmonics, and $\xi_{s}^{\sigma}$ are irreducible representations from the decomposition of the direct product of the spin states $\chi_{J_{1,a}}^{M_{1,a}}$ and $\chi_{J_{2,a}}^{M_{2,a}}$ of the individual particles,
@@
where the term with angle brackets is a Clebsch-Gordan coefficient. The sum over the total spin quantum number $s$ ranges from $\lvert J_{1,a}-J_{2,a}\rvert$ to $J_{1,a}+J_{2,a}$, and for each $s$ the sum over the third component $\sigma$ ranges from s-s to +s+s. The sum over the orbital-angular-momentum quantum number $l$ ranges from 0 to infinity, and for each $l$ the sum over the third component $m$ ranges from l-l to +l+l. The radial wavefunctions in Equation (18) are labeled by quantum numbers $\sigma$ and $m$ for the third components of the total spin and orbital angular momentum, respectively. It is more convenient to rewrite Equation (18) as an expansion in radial wavefunctions labeled by quantum numbers $J$ and $M$ for the total angular momentum,
@@
Once the above procedure has been applied to every channel $a$, one obtains infinitely many systems of radial Schrödinger equations labeled by conserved quantum numbers $J$ and $M$ for the total angular momentum and its third component. Using matrix notation, the system for given quantum numbers $J$ and $M$ can be written concisely as
$\biggl[\frac{1}{2\mu}\biggl(-\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}-\frac{2}{r}\frac{\mathrm{d}\hphantom{r}}{\mathrm{d}r}+\frac{L(L+1)}{r^{2}}\biggr)+V^{J}(r)-E\biggr]\psi^{JM}(r)=0,$ (21)
@@
where $L$ is a diagonal orbital-angular-momentum matrix, $V^{J}(r)$ is a radial potential matrix, and $\psi^{JM}(r)$ is a vector wavefunction. Note that the vector wavefunction depends on both $J$ and $M$, but the potential matrix depends only on $J$. The radial Schrödinger equations in this system are labeled by triplets $(a,l,s)$ that identify spin-orbital channels of the scattering process. Since $s$ and $l$ are constrained by the triangle relations $\lvert J_{1,a}-J_{2,a}\rvert\leq s\leq J_{1,a}+J_{2,a}$ and $\lvert J-s\rvert\leq l\leq J+s$, the system in Equation (21) contains only a finite number of radial Schrödinger equations for any value of $J$. To simplify the notation, we replace the labels $(a,l,s)$ for the spin-orbital channels by a single integer $i$. Then the radial potential matrix elements can be expressed as
@@
$V_{ij}^{J}(r)=\sum_{\sigma_{i},m_{i}}\sum_{\sigma_{j},m_{j}}\langle s_{i}l_{i}\sigma_{i}m_{i}|s_{i}l_{i}JM\rangle\langle s_{j}l_{j}\sigma_{j}m_{j}|s_{j}l_{j}JM\rangle\\ \times\int\mathrm{d}\Omega\,Y_{l_{i}}^{m_{i}}(\theta,\phi)^{\ast}Y_{l_{j}}^{m_{j}}(\theta,\phi)\xi_{s_{i}}^{\sigma_{i}{\dagger}}V_{a_{i}b_{j}}(\bm{r})\xi_{s_{j}}^{\sigma_{j}},$ (22)
@@
where $i\to(a_{i},s_{i},l_{i})$ and $j\to(b_{j},s_{j},l_{j})$. Note that each potential $V_{a_{i}b_{j}}(\bm{r})$ is an operator mapping spin vectors in the channel $b_{j}$ to spin vectors in the channel $a_{i}$ and that the total spin states $\xi_{s_{i}}^{\sigma_{i}}$ and $\xi_{s_{j}}^{\sigma_{j}}$ can be expanded in terms of the spin states of the individual particles in the channels $a_{i}$ and $b_{j}$ using Equation (19). Then we see that the right side of Equation (22) does not depend on $M$ by virtue of the Wigner-Eckart theorem.
@@
There is a somewhat common misunderstanding that the reduction of a three-dimensional Schrödinger equation to a radial form can be performed only for a spherically-symmetric potential that does not depend on the angles $\theta$ and $\phi$. However, in our derivation of Equations (21) and (22) we have made no assumption on the potential $V(\bm{r})$ other than it conserves angular momentum. In the trivial case where the potential is independent of spin, then conservation of angular momentum does indeed require that it is also spherically symmetric. If the potential does depend on spin, then it can also depend on the angles $\theta$ and $\phi$ but Equations (21) and (22) would still be valid. A brute-force method to determine if any given potential $V(\bm{r})$ conserves angular momentum is to evaluate the right side of Equation (22) with $J$ and $M$ in either one of the two Clebsch-Gordan coefficients replaced by $J^{\prime}$ and $M^{\prime}$. The potential $V(\bm{r})$ conserves angular momentum if and only if the result is zero for $J^{\prime}\neq J$ or $M^{\prime}\neq M$ and it is independent of $M$ when $J^{\prime}=J$ and $M^{\prime}=M$.
@@
The system in Equation (21) describes a scattering process taking place in a partial wave labeled by quantum numbers $J$ and $M$. Because of conservation of angular momentum, each partial wave is decoupled from the others. SPARSE solves the scattering problem only for one individual partial wave, leaving the task of summing partial waves to the user. Hence, from now on we will drop all superscripts for the conserved quantum numbers $J$ and $M$.
@@
To conclude this section, we simplify the derivatives inside Equation (21) as follows. Let us observe that